OpenStudy (anonymous):

f(x)=x^2+8x find f(x+h)-f(x)/h how would i solve this?

OpenStudy (ranga):

f(x) = x^2 + 8x f(x+h) = (x+h)^2 + 8(x+h) = x^2 + 2xh + h^2 + 8x + 8h = x^2 + (2h + 8)x + (h^2 + 8h) Subtract f(x) from f(x+h). Divide by h. Simplify.

OpenStudy (anonymous):

thanks professor! :D

OpenStudy (ranga):

You are welcome.